OpenStudy (anonymous):

alain says the ratios 3:5 and 2:10 are equal is he correct? explain. please help

OpenStudy (anonymous):

no. 3/5 isnt the same as 2/10

OpenStudy (mokeira):

not true 3:5 is same as \[\frac{ 3 }{ 5 }\] where as 2:10 is same as \[\frac{ 2 }{ 10 }=\frac{ 1 }{ 5 }\]